# Coherent polarization self-rotation

### v1

@@ Abstract
We introduce and study coherent polarization self-rotation (CPSR), a two-photon light-matter interaction in dense alkali-metal vapors that enables both narrowband optical spectroscopy of magnetic transitions and coherent coupling between light and collective atomic spins. Unlike conventional polarization self-rotation, CPSR requires initial spin polarization and a predominantly linearly polarized probe. It operates efficiently even in optically thick vapors with high buffer-gas pressure, rapid spin-exchange collisions, and optically-unresolved hyperfine structure. We demonstrate CPSR with near-unity contrast in rubidium and achieve an exceptionally narrow two-photon linewidth of 10 Hz in potassium. CPSR realizes a coherent interface between one optical quadrature and the long-lived collective electronic spin, offering a robust and scalable light-spin coupling in optically thick platforms. This opens new opportunities for quantum optics, including coherent transduction between light and ultra-long-lived noble-gas spins via alkali spins.
@@ I Introduction
Warm alkali-metal vapors provide a powerful platform for achieving both. They exhibit strong optical transitions at accessible wavelengths and reach high atomic densities with moderate heating ($\,{}^{\circ}$). Crucially, at elevated densities, these vapors can enter the spin-exchange-relaxation-free (SERF) regime, in which the Larmor coherence of the collective spin remains robust despite rapid interatomic spin-exchange collisions [9, 10]. Combined with their technological maturity and modern, compact vapor-cell designs, these features make alkali vapors ideal for high-performance and miniaturized spin-based devices [11, 12].
In practice, operating alkali vapor cells at elevated temperatures requires the use of buffer gases to suppress relaxation due to wall collisions. However, this homogeneously broadens the optical transitions, making the hyperfine structure optically unresolved [13, 14, 15, 16]. This broadening undermines most multi-photon spectroscopic techniques, such as electromagnetically induced transparency [17] and nonlinear magneto-optical rotation [18], which rely on resolving the hyperfine transitions. As a result, these techniques are restricted to low or moderate atomic densities and temperatures, making them incompatible with optimal spin-photon interfaces. Furthermore, when alkali spins are used to mediate coupling to the nuclear spins of noble gases, new schemes are required that remain effective in such broadened, optically-thick ensembles [19, 20, 21, 22].
We introduce and experimentally demonstrate a two-photon spectroscopic method—coherent polarization self-rotation (CPSR)—that overcomes these limitations. CPSR addresses the magnetic (Larmor) coherence of the collective electronic spin in alkali vapors. It operates effectively in dense, optically thick media with high buffer gas pressure, broadened optical lines, and fully unresolved hyperfine structure. It remains functional deep in the SERF regime and is immune to spin-exchange relaxation. The mechanism combines optical Faraday rotation and vector lightshift. Unlike standard polarization self-rotation (PSR) [23, 24, 25], CPSR relies on externally prepared spin polarization and a predominantly linearly polarized probe. We demonstrate CPSR spectroscopic lines with near-unity contrast in rubidium, indicating efficient light-to-spin mapping, and ultra-narrow 10 Hz-wide lines in potassium vapor. CPSR establishes a platform for coherent light-matter interfaces that remain effective in high-density, high-temperature ensembles, enabling strong coupling to long-lived collective spins and opening new directions in quantum sensing, optics, and hybrid systems involving noble-gas spins.
@@
II Light-spin coupling mechanism
We consider an alkali-vapor spin ensemble characterized by the collective hyperfine spin operators $\hat{\mathbf{F}}$. These are defined as $\hat{F}_{\mu}=\sum_{j=1}^{N_{\mathrm{a}}}\hat{F}^{j}_{\mu}$, where $\hat{F}^{j}_{\mu=\{x,y,z\}}$ are the hyperfine spin operators of the $j^{\text{th}}$ atom, and $N_{\mathrm{a}}$ is the number of atoms within the probe beam. Frequent spin-exchange collisions lead the spin populations to follow a spin-temperature distribution, such that $\hat{\mathbf{F}}$ effectively follows the electron spin, scaled by a factor $q$ known as the slowing-down factor [26, 27, 28]. This reduces the multilevel ground-state manifold to an effective spin-1/2 system (see Fig. 1a). We further assume that the spins are optically polarized along the $z$ axis (Fig. 1b, left), such that $\langle\hat{F}_{z}\rangle\gg\langle\hat{F}_{x}\rangle,\langle\hat{F}_{y}\rangle$, and $\langle\hat{F}_{z}\rangle=qN_{\mathrm{a}}p_{\mathrm{a}}/2$, where $-1\leq p_{\mathrm{a}}\leq 1$ is the mean spin polarization [29, 14]. An external magnetic field applied along $z$ induces Larmor precession at a frequency $\omega_{\mathrm{a}}$, at the kHz range or below.
@@ II Light-spin coupling mechanism
We present in Appendix B both a detailed theoretical model for CPSR and a simplified analytic solution, derived under specific assumptions discussed therein. The simplified model captures the essential features of the light-spin dynamics and yields a set of coupled equations for the transverse spin,
@@
Two points are worth emphasizing. First, $\sigma(\Delta)$ represents the true absorption cross section, with $\Gamma$ incorporating the pressure broadening from the buffer gas. Second, the spin decoherence rate $\gamma$ includes a contribution $\gamma_{\mathrm{se}}$ from spin-exchange collisions. When the spin-exchange collision rate $R_{\mathrm{se}}$ exceeds the Larmor frequency $\omega_{\mathrm{a}}$, the system enters the SERF regime, in which $\gamma_{\mathrm{se}}$ drops below $R_{\mathrm{se}}$ and scales quadratically with $\omega_{\mathrm{a}}$ [9, 26]. Both our theoretical treatment and all experimental measurements in this work are carried out in the SERF regime. Full expressions for $\gamma$ and $\gamma_{\mathrm{se}}$ in this limit are provided in Table 1 in Appendix B.
@@
Here, we define the complex, unitless response amplitude $\mathcal{C}(\omega)$ and identify the spin-probe coupling rate $\Omega=k_{{\mathrm{L}\rightarrow\mathrm{a}}}k_{{\mathrm{a}\rightarrow\mathrm{L}}}$ (see Appendix B). Notably, $\Omega$ scales with both the number of $z$-polarized atoms $\langle\hat{F}_{z}\rangle$ and the input photon flux $S_{1}^{\text{in}}$.
@@ III Rubidium CPSR spectroscopy
The atomic spins are subject to a magnetic field $B_{z}$ along the $z$-axis. They are optically pumped and held polarized along the $z$ direction using an auxiliary, circularly polarized $795\,\mathrm{nm}$ laser beam. The pumping beam, with power up to $100\,\mathrm{mW}$, covers the entire cell and is blue-detuned by 45 GHz from to the D1 transition. We estimate the corresponding pumping rate to be $R_{\mathrm{pump}}/q=23$ Hz. The spin decoherence rate is measured in the absence of optical pumping at low Larmor frequency, $\gamma(R_{\mathrm{pump}}=\omega_{\mathrm{a}}=0)=21.5$ Hz. This results in a spin polarization of $p_{\mathrm{a}}=0.52$ and a slowing-down factor of $q=7.4$ (see Appendices A and B).
@@
To demonstrate CPSR, we apply a magnetic field of $B_{z}\approx 1$ mG, corresponding to a Larmor frequency of $\omega_{\mathrm{a}}=268$ Hz including the pump-induced lightshift, and measure the transmission spectrum by scanning the polarization modulation frequency $\omega$. The amplitude and phase of the outgoing quadrature $S_{2}^{\text{out}}$ referenced to those of the probe beam before the cell are plotted in Fig. 2. The spectra exhibit on-resonance absorption and anti-resonance amplification near $\omega=\pm\omega_{\mathrm{a}}$, with a FWHM $\approx 2\gamma$ of approximately 100 Hz. The observed contrast reaches $|\mathcal{C}|=1.1$, exceeding unity, from which we infer a spin-light coupling rate of $\Omega=2\gamma|\mathcal{C}|\approx 110\,\mathrm{Hz}$. The near-complete extinction at $\omega=\omega_{\mathrm{a}}$ highlights the potential for highly efficient, coherent mapping of the light field onto the collective atomic spin.
@@
Numerical calculations based on our detailed theoretical model show excellent agreement with the experimental data, as demonstrated in Fig. 2. These calculations use independently measured physical parameters, with the exception of $n_{\mathrm{a}}$, $\gamma$, $\omega_{\mathrm{a}}$, and $\Delta$, which are fine-tuned within the experimental uncertainty (see Appendix B for details on the model and parameter values). The analytic solution of the simplified model (not shown) is quantitatively consistent with the results of the full numerical model.
@@
Figure 2: A complex spectrum of coherent polarization self-rotation. Measured transmission (a) amplitude and (b) phase of the optical signal traversing a rubidium vapor cell. The signal exhibits absorption at positive detuning and amplification at negative detuning, consistent with the predicted response near $\omega=\pm\omega_{\mathrm{a}}$. These spectral features, with a contrast exceeding unity and a full width below 100 Hz, arise from the dispersive Faraday interaction between the polarized collective alkali-spin and the polarization-modulated probe. The numerical solution of the detailed theoretical model (solid line) agrees well with the experimental data (circles).
@@
CPSR is inherently resistant to spin-exchange relaxation in the SERF regime. To demonstrate this, we measure transmission spectra across a range of Larmor frequencies satisfying $|\omega_{\mathrm{a}}|\ll R_{\mathrm{se}}$ by varying $B_{z}$. The results are shown in Fig. 3a. As expected, the CPSR spectral lines are significantly narrower than the bare spin-exchange relaxation rate and become progressively narrower as $|\omega_{\mathrm{a}}|$ decreases. Figures 3b and 3c summarize the extracted linewidth and contrast, respectively, as functions of $\omega_{\mathrm{a}}$. We observe a clear quadratic dependence of the linewidth on $\omega_{\mathrm{a}}$, characteristic of SERF behavior, along with a reduction of line contrast as the relaxation rate increases.
@@
Figure 3: Resistance of CPSR to spin-exchange relaxation. a. Transmission spectra of the optical signal for a range of Larmor frequencies $\omega_{\mathrm{a}}$. Dashed lines indicate the fitted linewidths (FWHM) of the absorption (blue) and amplification (red) peaks. The spectral narrowing at low $\omega_{\mathrm{a}}$ is a hallmark of the SERF regime, where Zeeman coherence is preserved despite frequent spin-exchange collisions. Black arrows indicate the data shown in Fig. 2. b. Dependence of the fitted FWHM on $\omega_{\mathrm{a}}$, along with the theoretical prediction (black dotted line), showing the expected quadratic dependence characteristic of the SERF mechanism. c. Contrast of the measured optical lines, extracted from the transmission extrema. Contrast exceeding unity is observed for the amplification peaks at low $\omega$. Dotted black lines in panels (b) and (c) are predictions of the detailed numerical solution.
@@
These results demonstrate the feasibility of achieving high-contrast, narrow optical spectral lines in optically thick alkali vapor under high buffer-gas pressure, conditions traditionally considered challenging for coherent light-spin coupling.
@@ IV Potassium CPSR spectroscopy with a 10-Hz linewidth
At a temperature of 185 C, the potassium vapor density was $n_{\mathrm{a}}=7.7\cdot 10^{13}\,\mathrm{cm^{-3}}$, yielding an on-resonance optical depth of $d(0)\approx 50$. The atoms were optically pumped using circularly polarized light at $770\,\mathrm{nm}$ with a power of $13\,\mathrm{mW}$, blue-detuned by $100\,\mathrm{GHz}$ from the atomic transition. A pump beam diameter of $4\,\mathrm{cm}$ ensured uniform pumping of the cell volume, reducing broadening of the CPSR lines. The probe beam, red-detuned by $215\,\mathrm{GHz}$ from the same transition, had a power of $25\,\mathrm{mW}$. For a Larmor frequency of $\omega_{\mathrm{a}}=29\,\mathrm{Hz}$ (magnetic field of $B_{z}=60~{}\mu$G), the measured potassium spin coherence time was 50 ms.
The corresponding CPSR spectrum, shown in Fig. 4, exhibits an exceptionally narrow optical linewidth of $9.7\,\mathrm{Hz}$, with a contrast of 20%. Notably, the resonance lineshapes deviate from the generalized Lorentzian form predicted by Eq. (4) due to optical pumping induced by the $S_{3}$ component of the probe itself. While this effect is neglected in the simplified model [Eqs. (1)], it is fully captured by the detailed model presented in Appendix B, which fits the data well.
Figure 4: CPSR in a spin-polarized potassium ensemble. The transmission spectrum demonstrates CPSR resonances at ±$\pm\omega_{\mathrm{a}}=\pm 29$ Hz with a full width of $9.72\gamma=$ 9.7 Hz and $\sim$20contrast. The solid lines are predictions of the numerical model (see Appendix B).

### v2

@@ Abstract
We introduce and study coherent polarization self-rotation (CPSR), a two-photon light-matter interaction in dense alkali-metal vapors that enables both narrowband optical spectroscopy of magnetic transitions and coherent coupling between light and collective atomic spins. Unlike conventional polarization self-rotation, CPSR requires initial spin polarization and a predominantly linearly polarized probe. It operates efficiently even in optically thick vapors with high buffer-gas pressure, rapid spin-exchange collisions, and optically-unresolved hyperfine structure. We demonstrate CPSR with near-unity contrast in rubidium and achieve an exceptionally narrow two-photon linewidth of 10 Hz in potassium. CPSR realizes a coherent interface between one optical quadrature and the long-lived collective electronic spin, offering a robust and scalable spin-light coupling in optically thick platforms. This opens new opportunities for quantum optics, including quantum-enhanced metrology in the audio-frequency band and coherent transduction between light and ultra-long-lived noble-gas spins via alkali spins.
@@ I Introduction
[…]
Crucially, at elevated densities, warm alkali-metal vapors can enter the spin-exchange-relaxation-free (SERF) regime [11, 12]. In this regime, the spin-exchange collision rate exceeds the Larmor frequency, locking the hyperfine and electron spin orientations to each other. As the total spin orientation is conserved during spin exchange collisions, the coherence time of the hyperfine spin orientation is extended. Its residual relaxation is caused by mixing of higher hyperfine moments, which still relax, due to non-vanishing Larmor precession. The SERF regime is realized at high temperatures and low magnetic fields, and allows for long, tens of milliseconds, coherence times at high optical depths.
In practice, operating alkali vapor cells at elevated temperatures requires the use of buffer gases to suppress relaxation due to wall collisions. However, this homogeneously broadens the optical transitions, making the hyperfine structure optically unresolved [13, 14, 15, 16]. This broadening undermines most multi-photon spectroscopic techniques, such as electromagnetically induced transparency [17] and nonlinear magneto-optical rotation [18]. Furthermore, when alkali spins are used to mediate coupling to the nuclear spins of noble gases through their spin orientations, new spin-light interfacing schemes are required that couple to alkali spin orientation [19, 20, 21, 22].
We introduce and experimentally demonstrate a two-photon spectroscopic method—coherent polarization self-rotation (CPSR)—that overcomes these limitations. CPSR addresses the magnetic (Larmor) coherence of the collective spin in alkali vapors. It operates effectively in dense, optically thick media with high buffer gas pressure, broadened optical lines, and fully unresolved hyperfine structure. It addresses the spin orientation and therefore remains functional deep in the SERF regime and is immune to spin-exchange relaxation. The mechanism combines optical Faraday rotation and vector lightshift. Unlike standard polarization self-rotation (PSR) [23, 24, 25], CPSR relies on externally prepared spin polarization and a predominantly linearly polarized probe. We demonstrate CPSR spectroscopic lines with near-unity contrast in rubidium, indicating efficient light-to-spin mapping, and 10 Hz-wide lines in potassium vapor. CPSR establishes a platform for coherent light-matter interfaces that remain effective in high-density, high-temperature ensembles, enabling strong coupling to long-lived collective spin states and opening new directions in quantum sensing, optics, and hybrid systems involving noble-gas spins.
@@
II Spin-light coupling mechanism
We consider an alkali-vapor spin ensemble characterized by the collective hyperfine spin operators $\hat{\mathbf{F}}$. These are defined as $\hat{F}_{\mu}=\sum_{j=1}^{N_{\mathrm{a}}}\hat{F}^{j}_{\mu}$, where $\hat{F}^{j}_{\mu=\{x,y,z\}}$ are the hyperfine spin operators of the $j^{\text{th}}$ atom, and $N_{\mathrm{a}}$ is the number of atoms within the probe beam. Frequent spin-exchange collisions lead the spin populations to follow a spin-temperature distribution, such that $\hat{\mathbf{F}}$ effectively follows the electron spin, scaled by a factor $q$ known as the slowing-down factor [26, 27, 28]. This reduces the multilevel ground-state manifold to an effective spin-1/2 system (see Fig. 1a). We further assume that the spins are optically polarized along the $z$ axis (Fig. 1b, left), such that $\langle\hat{F}_{z}\rangle\gg\langle\hat{F}_{x}\rangle,\langle\hat{F}_{y}\rangle$, and $\langle\hat{F}_{z}\rangle=qN_{\mathrm{a}}p_{\mathrm{a}}/2$, where $1\leq p_{\mathrm{a}}\leq 1$ is the mean spin polarization [29, 14]. An external magnetic field applied along $z$ induces Larmor precession at a frequency $\omega_{\mathrm{a}}$, at the kHz range or below.
@@ II Spin-light coupling mechanism
We present in Appendix B both a detailed theoretical model for CPSR and a simplified analytic solution, derived under specific assumptions discussed therein. The simplified model captures the essential features of the spin-light dynamics and yields a set of coupled equations for the transverse spin,
@@
Two points are worth emphasizing. First, $\sigma(\Delta)$ represents the detuned absorption cross section, with $\Gamma$ incorporating the pressure broadening from the buffer gas. Second, the spin decoherence rate $\gamma$ includes a contribution $\gamma_{\mathrm{se}}$ from spin-exchange collisions. When the spin-exchange collision rate $R_{\mathrm{se}}$ exceeds the Larmor frequency $\omega_{\mathrm{a}}$, the system enters the SERF regime, in which $\gamma_{\mathrm{se}}$ drops below the spin-exchange collision rate $R_{\mathrm{se}}$, scaling quadratically with $\omega_{\mathrm{a}}$, while higher hyperfine spin moments (e.g., spin alignment) relax [11, 26]. Both our theoretical treatment and all experimental measurements in this work are carried out in the SERF regime. Full expressions for $\gamma$ and $\gamma_{\mathrm{se}}$ in this limit are provided in Table 1 in Appendix B.
@@
Here, we define the complex contrast $\mathcal{C}(\omega)$ as the ratio between the Stokes parameter induced by the atomic medium and the transmitted input, and identify the spin-probe coupling rate $\Omega=k_{{\mathrm{L}\rightarrow\mathrm{a}}}k_{{\mathrm{a}\rightarrow\mathrm{L}}}$ (see Appendix B). Notably, $\Omega$ scales with both the number of $z$-polarized atoms $\langle\hat{F}_{z}\rangle$ and the input photon flux $S_{1}^{\text{in}}$.
@@
At the two-photon resonance, the atoms precess on the Bloch sphere with the modulation frequency of the Stokes parameters and follow their Poincaré trajectory with a $\pi/2$ phase delay (see Fig. 1b). Consequently, the atomic-induced $S_{2}$ response (the atomic readout) is $\pi$ out of phase with the input, leading to their destructive interference, measured as $S_{2}$ absorption. At the anti-resonance, the precession frequency is opposite to the Stokes modulation, such that while $F_{x},S_{3}$ behave similarly, $F_{y}$ mirrors its phase, leading to in-phase input and atomically-induced $S_{2}$, hence their constructive interference, measured as $S_{2}$ gain. In fact, examination of $S_{1}$ up to second order in the interaction strength reveals that absorption (gain) in $S_{2}$ is accompanied by fine attenuation (amplification) of $S_{1}$, reflecting a coherent signal-to-control exchange.
@@ III Rubidium CPSR spectroscopy
The atomic spins are subject to a magnetic field $B_{z}$ along the $z$-axis. They are optically pumped and held polarized along the $z$ direction using an auxiliary, circularly polarized $795\,\mathrm{nm}$ laser beam. The pumping beam, with power up to $100\,\mathrm{mW}$, covers the entire cell and is blue-detuned by 45 GHz from the D1 transition. The large pump detuning is chosen to mitigate pump depletion while preserving high spin polarization. Low pump depletion helps homogenize the spin polarization and decoherence, which is significant for the spectroscopic measurement. We estimate the corresponding pumping rate to be $R_{\mathrm{pump}}/q=23$ Hz. The spin decoherence rate is measured in the absence of optical pumping at low Larmor frequency, $\gamma(R_{\mathrm{pump}}=\omega_{\mathrm{a}}=0)=21.5$ Hz. This results in a spin polarization of $p_{\mathrm{a}}=0.52$ and a slowing-down factor of $q=7.4$ (see Appendices A and B).
@@
To demonstrate CPSR, we apply a magnetic field of $B_{z}\approx 1$ mG, corresponding to a Larmor frequency of $\omega_{\mathrm{a}}=268$ Hz including the pump-induced lightshift, and measure the transmission spectrum by scanning the polarization modulation frequency $\omega$. The amplitude and phase of the outgoing quadrature $S_{2}^{\text{out}}$ referenced to those of the probe beam before the cell are plotted in Fig. 2. The spectra exhibit on-resonance absorption and anti-resonance amplification near $\omega=\pm\omega_{\mathrm{a}}$, with a FWHM $\approx 2\gamma$ of approximately 100 Hz, primarily limited by pumping-induced relaxation (see Appendix A). The observed contrast reaches $|\mathcal{C}|=1.1$, exceeding unity, from which we infer a spin-light coupling rate of $\Omega=2\gamma|\mathcal{C}|\approx 110\,\mathrm{Hz}$. The near-complete extinction at $\omega=\omega_{\mathrm{a}}$ highlights the potential for highly efficient, coherent mapping of the light field onto the collective atomic spin.
@@
Numerical calculations based on the simplified model from Eq. (4) as well as the detailed theoretical model show excellent agreement with the experimental data, as demonstrated in Fig. 2. These calculations use independently measured physical parameters, with the exception of $n_{\mathrm{a}}$, $\gamma$, $\omega_{\mathrm{a}}$, and $\Delta$, which are fine-tuned within the experimental uncertainty (see Appendix B for details on the model and parameter values).
@@
Figure 2: A complex spectrum of coherent polarization self-rotation. Measured transmission (a) amplitude and (b) phase of the optical signal traversing a rubidium vapor cell. The signal exhibits absorption at positive detuning and amplification at negative detuning, consistent with the predicted response near $\omega=\pm\omega_{\mathrm{a}}$. These spectral features, with a contrast exceeding unity and a full width below 100 Hz, arise from the dispersive Faraday interaction between the polarized collective alkali-spin and the polarization-modulated probe. The numerical solution of the detailed theoretical model (dotted black line) and the simplified model from Eq. (4) (solid purple line) both agree well with the experimental data (circles).
@@
CPSR couples to the spin orientation and is therefore inherently resistant to spin-exchange relaxation in the SERF regime. To demonstrate this, we measure transmission spectra across a range of Larmor frequencies satisfying $|\omega_{\mathrm{a}}|\ll R_{\mathrm{se}}$ by varying $B_{z}$. The results are shown in Fig. 3a. As expected, the CPSR spectral lines are significantly narrower than the spin-exchange collisions rate $R_{\mathrm{se}}$ and become progressively narrower as $|\omega_{\mathrm{a}}|$ decreases. Figures 3b and 3c summarize the extracted linewidth and contrast, respectively, as functions of $\omega_{\mathrm{a}}$. We observe a clear quadratic dependence of the linewidth on $\omega_{\mathrm{a}}$, characteristic of SERF behavior, along with a reduction of line contrast as the relaxation rate increases.
@@
Figure 3: Resistance of CPSR to spin-exchange relaxation. a. Transmission spectra of the optical signal for a range of Larmor frequencies $\omega_{\mathrm{a}}$. Dashed lines indicate the fitted linewidths (FWHM) of the absorption (blue) and amplification (red) peaks. The spectral narrowing at low $\omega_{\mathrm{a}}$ is a hallmark of the SERF regime, where Zeeman coherence is preserved despite frequent spin-exchange collisions. Black arrows indicate the data shown in Fig. 2. b. Dependence of the fitted FWHM on $\omega_{\mathrm{a}}$, along with the theoretical prediction (black dotted line), showing the expected quadratic dependence characteristic of the SERF mechanism. c. Contrast of the measured optical lines, extracted from the transmission extrema. Contrast exceeding unity is observed for the amplification peaks at low $\omega$. In panels (b) and (c), dotted black and solid purple lines are the predictions of the detailed numerical model and simplified model [Eq. (4)], respectively.
@@
These results demonstrate the feasibility of achieving high-contrast, narrow optical spectral lines in optically thick alkali vapor under high buffer-gas pressure, conditions traditionally considered challenging for coherent spin-light coupling.
@@ IV Potassium CPSR spectroscopy with a 10-Hz linewidth
[…]
The corresponding CPSR spectrum, shown in Fig. 4, exhibits a narrow optical linewidth of $9.7\,\mathrm{Hz}$, with a contrast of 20%. Notably, the resonance lineshapes deviate from the generalized Lorentzian form predicted by Eq. (4) due to optical pumping induced by the $S_{3}$ component of the probe itself. The linewidth here is limited by this probe-induced pumping, as well as by depopulation following potassium-helium collisions. While this effect is neglected in the simplified model (Eqs. (1) ), it is fully captured by the detailed model presented in Appendix B, which fits the data well.
Figure 4: CPSR in a spin-polarized potassium ensemble. The transmission spectrum demonstrates CPSR resonances at $\pm\omega_{\mathrm{a}}=\pm 29$ Hz with a full width of $2\gamma=9.7$ Hz and $\sim$20% contrast. The data (blue circles) deviates from the simplified model’s prediction [Eq. (4), solid purple lines] due to non-negligible optical pumping by the input $S_{3}$ component, which is captured by the numerical model (dotted black lines)
(see Appendix B).